OpenStudy (anonymous):

Use the fundamental identities to simplify this expression :csc x/cotxsecx PLEASE SHOW ME STEP BY STEP HOW TO GET THE ANSWER!

OpenStudy (anonymous):

Well, let's split this up so we can make this more visual then: \[\frac{ cscx }{ cotx secx} = \frac{ cscx }{ 1 }\cdot \frac{ 1 }{ cotx } \cdot \frac{ 1 }{ secx }\]So what would all of these expressions look like as sines and cosines?

OpenStudy (anonymous):

Wait which identity did you use?

OpenStudy (anonymous):

No identity at all. I simply split up the multiplication to allow us to look at each individual piece more easily. I did the same thing as if I were to say: \[\frac{ 2 }{ 3 } = \frac{ 2 }{ 1 }\cdot \frac{ 1 }{ 3 }\] No identity at all.

OpenStudy (anonymous):

OH ok i got ya, what do you mean look like sins and cosines

OpenStudy (anonymous):

Well, how would I turn cscx into something involving sinx? Same for the other two, if I have 1/cotx, what would that be if I wanted to make it into sines and cosines, etc.

OpenStudy (anonymous):

well for cotx you could do cosx/sinx. For secx i could change i to 1/cosx

OpenStudy (anonymous):

it